OpenStudy (anonymous):

change e^2x=75 to logarithmic form

OpenStudy (anonymous):

ln(75)/2 = x

OpenStudy (earthcitizen):

\[e ^{2x}=75\] \[2x \ln e=\ln 75\] \[2x=\ln 75 \therefore x=(\ln 75)/2\]

OpenStudy (earthcitizen):

x=2.159

OpenStudy (anonymous):

use the properties of logarithms to write the expression, log \[\sqrt[5]{3y/x}\] as a sum, difference, or multiples of logarithms of x, y, or constants. assume that x>0 and y>0

OpenStudy (anonymous):

(its log of that square root it wouldnt go together)